# Sorbonne Université & CNRS, 4 place Jussieu, 75252 Paris Cedex 05, France

### v1

@@ 4.2.5 Transformers
Transformer models, first introduced in vaswani2017attention , have revolutionised deep learning by enabling highly parallel processing and effectively capturing long-range dependencies. In particular, their application in high-energy physics have already shown promising results in analysing complex particle interactions Qu:2022mxj ; Wu:2024thh ; Hammad:2023sbd ; Hammad:2024cae . Unlike other neural networks, transformers rely on a multi-head self-attention mechanism that enables global context modelling across input elements. This makes them particularly suitable for studying the intricate and highly correlated structure of particle clouds Qu:2019gqs in collider events and assess the contribution of each particle to the overall event structure to allow for improved signal and background classification.
@@
The transformer architecture employed in this work is illustrated in figure 9 and begins with a series of fully connected embedding layers consisting of 512, 256 and 128 neurons. The resulting feature representation, of dimension $(9,128)$, is passed through a normalisation layer and subsequently through a stack of 10 self-attention blocks. Each block contains 16 attention heads and incorporates a residual connection followed by Root Mean Square Layer Normalisation (RMSNorm) zhang2019root . The output is then processed by two fully connected layers with Sigmoid-Weighted Linear Unit (SiLU) activations elfwing2018sigmoid . After the final self-attention block, average pooling is applied across the particle dimension. The pooled vector is fed into a classification head composed of two dense layers with 128 and 64 units, respectively. Training is conducted over 21 epochs using a batch size of 128 and binary cross entropy as the loss function. Optimisation is performed using the AdamW algorithm with an initial learning rate of $0.0005$, adjusted during training via a cosine annealing scheduler loshchilov2016sgdr .
@@ 5 Results
For each of the three network architectures and each of the three FCNC signal processes on which they have been trained (yielding a total of nine trained models), the number of signal events $N_{s}$ passing the final selection is defined as the sum of the contributions from single production (SP) and pair production (PP). Considering one flavour scenario at a time, we thus define
@@
Additionally, the expected number of background events is determined independently for each of the nine models, based on the corresponding trained network and the optimised signal selection. Exclusion limits are then derived using the Pyhf library Heinrich:2021gyp , employing the $\tilde{q}_{\mu}$ test statistic Cowan:2010js with a luminosity-dependent relative uncertainty for the background yield. The latter is fixed to be 10% at $\mathcal{L}=100~{}\text{fb}^{-1}$ and then scaled with luminosity according to Poisson statistics,
@@
For each network architecture, the final ${\rm CL}_{s}$ value is taken to be the highest among the different signal-specific models, thus conservatively ignoring the potential gain that could stem from the combination of our three signal regions, which is hard to achieve without a proper assessment of the uncertainty correlations between the signal regions.
@@
Figure 11 presents the expected 95% confidence level (CL) exclusions on the anomalous couplings $f_{tu}^{\gamma}/\Lambda$ (top row) and $f_{tc}^{\gamma}/\Lambda$ (bottom row) as a function of the integrated luminosity for each of the three considered networks, namely our MLP (left), GAT (centre) and transformer (right) architectures. As anticipated, the bounds on FCNC couplings involving the up quark are stronger than those involving the charm quark. This trend is visible across all models, with the MLP yielding an improvement by a factor of approximately 1.6 for $f_{tu}^{\gamma}$ over $f_{tc}^{\gamma}$ while the GAT and transformer models show a milder factor of about $1.2-1.3$. This discrepancy is primarily driven by the kinematic and partonic structure of the LHC: the $u$-quark parton distribution function is significantly larger than that of the $c$-quark, leading to higher rates for the single top production mechanism mediated by $f_{tu}^{\gamma}$. Under the assumption $f_{tu}^{\gamma}=f_{tc}^{\gamma}$, the corresponding cross sections indeed differ by nearly a factor of 7 in favour of the $u$-quark channel, as shown in figure 2. However, this difference is partially offset by the better sensitivity to charm-quark FCNC in the pair production channel where the top decay $t\to c\gamma$ leads to a more distinctive final state than the $t\to u\gamma$ decay. This balance between production rate and classifier efficiency helps explain why the transformer and GAT models display a reduced gap between the two flavour scenarios compared to the MLP.
@@ 6 Summary and conclusions
We began by evaluating the sensitivity of a traditional cut-based analysis using a set of manually optimised kinematic selections. Although this strategy provides a useful baseline, its performance is ultimately limited by the inability to capture complex correlations among observables. The resulting signal significances were modest, underscoring the need for more advanced classification approaches. To address this, we employed three deep learning architectures to perform a supervised binary classification task designed to separate signal from background. Among these, MLPs exhibited the lowest performance, with an AUC of 0.928 for the single top channel and approximately 0.78 for top pair production, this resulting from solely relying on event-level kinematic variables. An analysis of feature importance showed that the classification performance was dominated by a small subset of inputs including the $H_{T}$, $\Delta R(b,\ell)$ and $p_{T}^{\gamma}$ observables, which together account for nearly 80% of the network’s predictive power.
@@
The improved classification capabilities of the GAT and transformer networks translated directly into more stringent bounds on the FCNC couplings. Assuming an integrated luminosity of $\mathcal{L}=100~{}{\rm fb}^{-1}$ and a $10\%$ background uncertainty, we found that the transformer and GAT models yield expected limits on the anomalous couplings $f_{tu}^{\gamma}/\Lambda$ and $f_{tc}^{\gamma}/\Lambda$ that are stronger by a factor of approximately $1.3-1.7$ compared to the MLP, that itself outperforms the performance of our simpler cut-based analysis. At $\mathcal{L}=139~{}{\rm fb}^{-1}$, values of $f_{tu}^{\gamma}/\Lambda$ down to $7.3-9.2\times 10^{-3}~{}{\rm TeV}^{-1}$ and $f_{tc}^{\gamma}/\Lambda$ down to $8.7-14.2\times 10^{-3}~{}{\rm TeV}^{-1}$ can hence be excluded, which will improve further at the HL-LHC where coupling values as small as $2\times 10^{-3}~{}{\rm TeV}^{-1}$ may be probed. The exclusions on the corresponding FCNC branching ratios follow a similar trend. At $\mathcal{L}=139~{}{\rm fb}^{-1}$, we find that the MLP excludes ${\rm BR}(t\to u\gamma)$ and ${\rm BR}(t\to c\gamma)$ down to about $1.0\times 10^{-5}$ and $2.4\times 10^{-5}$, respectively, with the GAT and transformer architectures improving these bounds by factors of $1.5-2.6$. At the HL-LHC, branching fractions as small as $10^{-6}$ could eventually be excluded at the 95% confidence level using transformer-based classification.
In summary, our results demonstrate the power of incorporating deep learning approaches into collider analyses, particularly for rare top quark decay processes. With the vast increase in statistics expected from future LHC runs, traditional cut-based methods will become increasingly suboptimal. Our findings suggest that attention-based architectures, such as transformers and GATs, offer significant potential for improving the sensitivity to new physics signatures in top quark processes at the LHC and beyond.

### v2

@@ 4.2.5 Transformers
Transformer models, first introduced in Vaswani et al. (2017a), have revolutionised deep learning by enabling highly parallel processing and effectively capturing long-range dependencies. In particular, their application in high-energy physics have already shown promising results in analysing complex particle interactions Qu et al. (2022); Wu et al. (2025); Hammad et al. (2024b); Hammad and Nojiri (2024). Unlike other neural networks, transformers rely on a multi-head self-attention mechanism that enables global context modelling across input elements. This makes them particularly suitable for studying the intricate and highly correlated structure of particle clouds Qu and Gouskos (2020) in collider events and assess the contribution of each particle to the overall event structure to allow for improved signal and background classification.
@@
The transformer architecture employed in this work is illustrated in figure 9 and begins with a series of fully connected embedding layers consisting of 512, 256 and 128 neurons. The resulting feature representation, of dimension $(9,128)$, is passed through a normalisation layer and subsequently through a stack of 10 self-attention blocks. Each block contains 16 attention heads and incorporates a residual connection followed by Root Mean Square Layer Normalisation (RMSNorm) Zhang and Sennrich (2019). The output is then processed by two fully connected layers with Sigmoid-Weighted Linear Unit (SiLU) activations Elfwing et al. (2018). After the final self-attention block, average pooling is applied across the particle dimension. The pooled vector is fed into a classification head composed of two dense layers with 128 and 64 units, respectively. Training is conducted over 21 epochs using a batch size of 128 and binary cross entropy as the loss function. Optimisation is performed using the AdamW algorithm with an initial learning rate of $0.0005$, adjusted during training via a cosine annealing scheduler Loshchilov and Hutter (2016).
@@ 5 Results
For each of the three network architectures and each of the three FCNC signal processes on which they have been trained (yielding a total of nine trained models), the number of signal events $N_{s}$ passing the final selection is defined as the sum of the contributions from single production (SP) and pair production (PP). In the training procedures, background events are always combined according to their weight (shown in Table 1) and the considered events account for the preselection cuts. Considering one flavour scenario at a time, we thus define
@@
Additionally, the expected number of background events is determined independently for each of the nine models, based on the corresponding trained network and the optimised signal selection. Exclusion limits are then derived using the Pyhf library Heinrich et al. (2021), employing the $\tilde{q}_{\mu}$ test statistic Cowan et al. (2011) with a luminosity-dependent relative uncertainty for the background yield. The latter is fixed to be 10% at $\mathcal{L}=100~\text{fb}^{-1}$ and then scaled with luminosity according to Poisson statistics,
@@
For each network architecture, the final ${\rm CL}_{s}$ value is taken to be the highest among the different signal-specific models when applied to the full signal combining the single and pair production channels, thus conservatively ignoring the potential gain that could stem from the combination of our three signal regions, which is hard to achieve without a proper assessment of the uncertainty correlations between the signal regions.
@@
Figure 11 presents the expected 95% confidence level (CL) exclusions on the anomalous couplings $f_{tu}^{\gamma}/\Lambda$ (top row) and $f_{tc}^{\gamma}/\Lambda$ (bottom row) as a function of the integrated luminosity for each of the three considered networks, namely our MLP (left), GAT (centre) and transformer (right) architectures. As anticipated, the bounds on FCNC couplings involving the up quark are stronger than those involving the charm quark. This trend is visible across all models, with the MLP yielding an improvement by a factor of approximately 1.6 for $f_{tu}^{\gamma}$ over $f_{tc}^{\gamma}$ while the GAT and transformer models show a milder factor of about 1.21.31.2-1.3. This discrepancy is primarily driven by the kinematic and partonic structure of the LHC: the $u$-quark parton distribution function is significantly larger than that of the $c$-quark, leading to higher rates for the single top production mechanism mediated by $f_{tu}^{\gamma}$. Under the assumption $f_{tu}^{\gamma}=f_{tc}^{\gamma}$, the corresponding cross sections indeed differ by nearly a factor of 7 in favour of the $u$-quark channel, as shown in figure 2. However, this difference is partially offset by the better sensitivity to charm-quark FCNC in the pair production channel where the top decay $t\to c\gamma$ leads to a more distinctive final state than the $t\to u\gamma$ decay. This balance between production rate and classifier efficiency helps explain why the transformer and GAT models display a reduced gap between the two flavour scenarios compared to the MLP.
@@ 6 Summary and conclusions
We began by evaluating the sensitivity of a traditional cut-based analysis to obtain baseline results, the performance being obviously limited by the analysis inability to capture complex correlations among observables. To address this, we next employed three deep learning architectures to perform a supervised binary classification task designed to separate signal from background. Among these, MLPs exhibited the lowest performance, with an AUC of 0.928 for the single top channel and approximately 0.78 for top pair production, this resulting from solely relying on event-level kinematic variables. An analysis of feature importance showed that the classification performance was dominated by a small subset of inputs including the $H_{T}$, $\Delta R(b,\ell)$ and $p_{T}^{\gamma}$ observables, which together account for nearly 80% of the network’s predictive power.
@@
The improved classification capabilities of the GAT and transformer networks translated directly into more stringent bounds on the FCNC couplings. Assuming an integrated luminosity of $\mathcal{L}=100~{\rm fb}^{-1}$ and a $10\%$ background uncertainty, we found that the transformer and GAT models yield expected limits on the anomalous couplings $f_{tu}^{\gamma}/\Lambda$ and $f_{tc}^{\gamma}/\Lambda$ that are stronger by a factor of approximately 1.31.71.3-1.7 compared to the MLP. At $\mathcal{L}=139~{\rm fb}^{-1}$, values of $f_{tu}^{\gamma}/\Lambda$ down to $7.3-9.2\times 10^{-3}~{\rm TeV}^{-1}$ and $f_{tc}^{\gamma}/\Lambda$ down to $8.7-14.2\times 10^{-3}~{\rm TeV}^{-1}$ can hence be excluded, which will improve further at the HL-LHC where coupling values as small as $2\times 10^{-3}~{\rm TeV}^{-1}$ may be probed. The exclusions on the corresponding FCNC branching ratios follow a similar trend. At $\mathcal{L}=139~{\rm fb}^{-1}$, we find that the MLP excludes ${\rm BR}(t\to u\gamma)$ and ${\rm BR}(t\to c\gamma)$ down to about $1.0\times 10^{-5}$ and $2.4\times 10^{-5}$, respectively, with the GAT and transformer architectures improving these bounds by factors of 1.52.61.5-2.6. At the HL-LHC, branching fractions as small as $10^{-6}$ could eventually be excluded at the 95% confidence level using transformer-based classification.
In summary, our results demonstrate that with the vast increase in statistics expected from future LHC runs, traditional machine-learning methods such as MLPs will become increasingly suboptimal. Our findings suggest that attention-based architectures, such as transformers and GATs, offer significant potential for improving the sensitivity to new physics signatures in top quark processes at the LHC and beyond.